Sailing to Miami. At point A the captain of a ship
determined that the distance to Miami was 13 miles.
If she sailed north to point B and then west to Miami,
the distance would be 17 miles. If the distance from
point A to point B is greater than the distance from
point B to Miami, then how far is it from point A to
point B?

Answer by mananth(16946)

The directions given indicate that point A Point B and Miami form a right angled triangle.
Point A too Miami becomes the hypotenuse which is 13 miles
Let distance from A to B be x
Then distance from B to Miami will be 17-x ( because AB + B to Miami is 12)
13^ = x^2+ (17-x)^2
169 = x^2+ (289-34x+x^2)
169 =x^2+289 -34x +x^2
2x^2-34x+ 120=0
x^2 -17x +60=0
x^2-12x-5x+60=0
x(x-12)-5(x-12)=0
(x-12)(x-5) =0
x=12 or 5